show that if (A/B)=1, then P(B^c/A^c)= 1
1
Expert's answer
2021-09-20T15:59:02-0400

Definition of complement

"P(B^C|A^C)=1 -P(B|A^C)"

Bayes' rule

"P(B^C|A^C) = 1 -(\frac{P(A^C|B)P(B)}{P(A^C)})"

Definition of complement

"P(B^C|A^C) = 1 -(\frac{(1 -P(A|B))P(B)}{P(A^C)})"

Using "P(A|B) = 1"

"P(B^C|A^C) = 1 -(\frac{(0)P(A)}{P(B^C)}) \\

P(B^C|A^C) = 1"